OpenStudy (anonymous):

What statement makes the open sentence 49 – 7x = 21 true? A. x = 11 B. x = 7 C. x = 4 D. x = 3

OpenStudy (lncognlto):

Well, rearrange the equation: 49 - 7x = 21, so -7x = 21 - 49, ao -7x = -28, so 7x = 28, so then x = 28/7. What is the value of x/

OpenStudy (lncognlto):

*so

OpenStudy (anonymous):

It is C. because 28/7=4